Problem 6 A website for home pregnancy test cites the following: When the subjects using the test were women who collected aClear and detailed answers appreciated. I will like lifesaver!
Answer #1

TOPIC:Application of the Bayes' theorem.

Prooblem - 6% Define the following events as L T : CP: The test is positive; The subject is pregnant. Given that, ay sensitivAgain, P (of being actually pregnant) = 30 = 0.30 XP (P) = 0.30. I Law of comple- mentation) -P(P) = 1-0.30 = 0.70. We need t